OpenStudy (crashonce):

A fair five sided dice with sides 1-5 is tossed 3 times. Find the probability that at least two will be odd

OpenStudy (crashonce):

@ganeshie8

OpenStudy (crashonce):

@Hero @kropot72

ganeshie8 (ganeshie8):

whats the probability for getting an odd number in a single toss ?

OpenStudy (crashonce):

3/5

ganeshie8 (ganeshie8):

good, you want to find probability for atleast two toss results being odd

ganeshie8 (ganeshie8):

that means you can have : 1) an odd number on exactly two tosses or 2) an odd number on all three tosses yes ?

OpenStudy (crashonce):

yes

ganeshie8 (ganeshie8):

find the probability of each and add up

ganeshie8 (ganeshie8):

whats the probability of having `an odd number on exactly two tosses` ?

OpenStudy (crashonce):

not sure can u show me

ganeshie8 (ganeshie8):

odd number on exactly two tosses means, one toss has to be even, yes ?

OpenStudy (crashonce):

yes

ganeshie8 (ganeshie8):

probability for odd number on a toss = 3/5 probability for even number on a toss = 2/5 lets count how many ways we can have an `odd number on exactly two tosses` : 1) (3/5)(3/5)(2/5) 2) (3/5)(2/5)(3/5) 3) (2/5)(3/5)(3/5)

ganeshie8 (ganeshie8):

Easy! just notice that each way has probabilties for "two odd numbers" and "one even number"

ganeshie8 (ganeshie8):

the even number could appear on third, second, or first toss ^

ganeshie8 (ganeshie8):

so the probability for `having odd number on exactly two tosses` would be : 3*(3/5)(3/5)(2/5) = ?

OpenStudy (crashonce):

36/125?

ganeshie8 (ganeshie8):

check again

OpenStudy (crashonce):

oh 54/125

ganeshie8 (ganeshie8):

Yes! next work the probability for having an odd number on ALL 3 dice (this should be easy)

OpenStudy (crashonce):

81/125

ganeshie8 (ganeshie8):

try again, step by step

OpenStudy (crashonce):

isn't it 3*(3/5 3/5 3/5)

ganeshie8 (ganeshie8):

probability for having an odd number on ALL 3 dice : (3/5) * (3/5) * (3/5) thats all

OpenStudy (crashonce):

why is there no 3

ganeshie8 (ganeshie8):

you don't need to multiply by 3 here because you dont have any cases here

ganeshie8 (ganeshie8):

earlier you had cases because the "even number" can appear on either first or second or third dice

ganeshie8 (ganeshie8):

working probability for an odd number on ALL 3 dice is pretty straightforward you know probability for odd number in a single toss = 3/5 since you want an odd number on ALL 3 dice, simply multiply the probabilties : (3/5)(3/5)(3/5)

OpenStudy (crashonce):

ok so is the final answer just the two added?

ganeshie8 (ganeshie8):

yes just add both the probabilities

OpenStudy (crashonce):

so its 81/125 thanks again ganeshie

ganeshie8 (ganeshie8):

looks good !